007 10.0 points A potters wheel moves from rest to an angu- lar speed of 0.10 rev/s in 26.9 s. Assuming constant angular acceleration, what is its angular acceleration in rad/s7 2 Answer in units of rad/s
Answer #1

By using,

w = wo + at

Where, wo = 0

w = 0.1 rev/s *(2*pi/ 1 rev) = 0.1*2*pi rad/s

a = (w - wo​​​​​​) /t

= (0.1*2*pi - 0)/26.9

= 0.023 rad/s^2